• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name zana

    Log entry time 23:53:53 on June18,2010

    Entry number 327443

    keyword=Shift Summary Swing shift

    16:00 Current at 70 uA : Higher temperature for Downstram collimator at 507C
    16:20 High voltage for Compton keep tripping with beam over 65uA. Alexandre and Megan set the Compton HV to 1700V. Now is stable. Megan is going to work with the MCC for solving the problem
    17:00 Temperature for the Dowstream collimator at 528C
    17:30 Megan did a spot++ at 5nA. The target surfice looks still uniform
    18:00 From MCC they are doing some tuning for the Compton
    18:55 As suggested by Bob Michaels, we did again a spot++, this time with the Slit at 15. We checked if we were hitting the target (as suggested by the RC), checking the T5 rates with beam ON and OFF (at 5nA)
    19:10 Target changed to Thick Carbon, in order to do the Compton tuning
    21:20 When just we wanted to give up on Compton, is possible they found a fix. Started fast beam recovery procedure. Problem with the beam position for BPMA X . Notified MCC: They have problem mantaining lower rates for the Compton and the BPMA X position at the same time
    22:00 MCC solved their problem with the beam position, we asked for 70uA
    22:30 Beam is at 70uA and Compton is runnning, checked Fast Feedback, Charge feedback and Dithering